OpenStudy (anonymous):

is the equation true, false, or open? 4y + 8 = 6y + 3? please i need help

OpenStudy (anonymous):

y = 5 / 2 = 2.5

OpenStudy (anonymous):

Welcome to Open Study

OpenStudy (anonymous):

so is it open or true

OpenStudy (austinl):

Please refrain from giving answers @LordHades Hello and welcome to OpenStudy @nickky9875 You want to get your terms with variables on one side and constants on the other.

OpenStudy (anonymous):

ok so i dont need the answer i need to know if it is false open or true its not for homework im just trying to prove it is a true statement is it?

OpenStudy (anonymous):

@austinL My apologies.

OpenStudy (anonymous):

please

OpenStudy (austinl):

Plug in a arbitrary value for y. Does the equality hold true?

OpenStudy (anonymous):

ummm yes

OpenStudy (austinl):

Put in y=1, what do you get?

OpenStudy (anonymous):

idk what do you get

OpenStudy (austinl):

\(4y+8=6y+3\) \(4(1)+8=6(1)+3\) \(12\ne9\) Any thoughts on this?

OpenStudy (anonymous):

yess so what is it

OpenStudy (anonymous):

is that you

OpenStudy (anonymous):

True